OpenStudy (anonymous):

Part 1: Solve the inequality 2(x + 6) greater than or equal to 3x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

jhonyy9 (jhonyy9):

2(x+6)>= 3x-2 2x+12>=3x-2 14>=x